A reaction has a rate constant of 7.3x10 M/s at 315 K and a frequency factor of 4.4x10 What is the activation energy of the reaction?

1
Expert's answer
2022-04-21T07:26:03-0400

Solution:

Activation energy of the reaction can be calculated from Arrhenius equation

k = A × e−(Ea/RT)

where,

k = chemical reaction rate

A = frequency factor

Ea = activation energy

R = gas constant

T = temperature


k = 7.3×10−4 M s−1

A = 4.4×107 M s−1

Ea = ???

R = 8.314 J mol−1 K−1

T = 315 K


Thus,

(7.3×10−4) = (4.4×107) × e−Ea / (8.314 × 315)

1.659×10−11 = e−Ea × 0.0003818

ln(1.659×10−11) = −Ea × 0.0003818

−24.8222 = −Ea × 0.0003818

Ea = 65013.67 J mol−1 = 6.5×104 J mol−1

Ea = 6.5×104 J/mol


Answer: The activation energy of the reaction is 6.5×104 J/mol